# Implications of computer science theory for the simulation hypothesis

### v2

@@ Abstract
The simulation hypothesis has recently excited renewed interest, especially in the physics and philosophy communities. However, the hypothesis specifically concerns computers that simulate physical universes, which means that to formally investigate it we need to couple computer science theory with physics. Here I couple those fields with the physical Church-Turing thesis. I then exploit that coupling to investigate of some of the computer science theory aspects of the simulation hypothesis. In particular, I use Kleene’s second recursion theorem to prove that it is mathematically possible for us to be a simulation that is being run on a computer — by us. In such a self-simulation, there would be two identical instances of us; the question of which of those is “really us” is meaningless. I also show how Rice’s theorem provides some interesting impossibility results concerning simulation and self-simulation; briefly describe the philosophical implications of fully homomorphic encryption for (self-)simulation; and briefly investigate the graphical structure of universes simulating universes simulating universes $\ldots$, among other issues. I end by describing some of the possible avenues for future research.
@@ 1.1 Background
The “simulation hypothesis” is an intriguing concept that in one or form or another has existed for thousands of years, in many human cultures. Since the advent of powerful digital computers over the last two decades, it has excited renewed interest among physicists as well as philosophers of science.
This modern iteration of the hypothesis has many versions that vary in their precise details. However, the central idea is that some portion of a physical universe, including some conscious reasoning agents that exist in that cosmological universe, might in fact be part of a simulation being run in a physical computer of some different, super-sophisticated race of aliens [12, 31, 20]. Under the simulation hypothesis, those “conscious reasoning agents” would just be variables evolving in a computer program running in that physical computer designed by that putative super-sophisticated race of aliens.
Much of philosophy of science, going back to Kant (at least), would be rendered moot if we are simulations. In particular, this would be the case for many of the flavors of “*** realism” (e.g., scientific realism), which suppose there “exists” some “real” physical truth which is not reducible to a mathematical formalism, a physical truth that instead is somehow “concrete”, and non-mathematical in its very essence — and so not just a computer algorithm. [^1] Arguably, the simple fact that the simulation hypothesis is logically consistent, but if true would imply that there is no such “concrete”, non-mathematical reality, establishes that it is impossible to establish the logical necessity of any of these flavors of realism,
The central idea of the simulation hypothesis has been extended in an obvious manner, simply by noting that the aliens that simulate our universe might themselves be a simulation in the computer of some even more sophisticated species, and so on and so on, in a sequence of ever-more sophisticated aliens. Similarly, going the other direction, in the not too distant future we might produce our own simulation of a universe running in some future computer that we will create, a simulation complete with variables that constitute “conscious, reasoning agents”. Indeed, we might produce such a simulation in which the reasoning agents can produce their own simulated universe in turn, etc. So in the near future, there might be a sequence of species’, each one with a computer running a simulation that produces the species just below it in the sequence, with us somewhere in the “middle” of that sequence.
The literature of the last few decades on the simulation hypothesis has focused almost exclusively on whether we, in our universe are such a simulation. This question is answered rather trivially if we adopt the view of ontic structural realism, especially if that view is formalized in terms of Tegmark’s level IV multiverse [59, 60]: yes, in some universes we are a simulation, and no, in some other universes we are not. It is not an either / or ’nuf said. (See also [16, 49].)
These investigations of relative probabilities all beg (many) questions, about what such probabilities might actually mean, formally. (Just as there are such questions concerning the idea of ascribing probabilities to the universes in a level IV multiverse in general.) How does one ascribe a measure, obeying the Kolmogorov axioms based on some associated sigma algebra, to a collection of events each of which is a universe? Note that it is not even clear that this collection would be a set, rather than a proper class of some sort. Directly reflecting such problems, one could not experimentally assess any proposed value of such a probability, e.g., with a proper scoring rule. [^2] One might try to use a Bayesian “degree of belief” interpretation of probability to circumvent this issue. However, without any decision that would be made based on the probability assigned, and associated loss function, one could not apply Bayesian decision theory. So it is hard to see how “degree of belief” is meaningful in this case. See [62] for some related arguments.
More prosaically, many papers considering such probabilities assume, implicitly or otherwise, that there is some way to assign a uniform probability to all universes. It is proven below in Section 2.4 that this is mathematically those in the kind of universes considered in this paper (which, one might argue, is the kind of universes considered in these other papers). One might imagine that rather than a uniform probability distribution, one could assign some sort of Cantor (“fair coin”) measure to the set of possible universes, as is done for example in algorithmic information theory. However, it is proven in Section 7 that it is impossible to assign a Cantor measure to a set of universes obeying a natural set of restrictions.
In contrast to this work on the probability that we are living in a simulation, some physicists have focused on whether there might be ways of experimentally determining whether our universe is a program in a simulation [13, 11, 9]. Work in this vein can only provide partial answers to the question of whether we are a simulation, at best. For example, some of this work first make the assumption that the simulating computer is digital, and so cannot fully faithfully represent real numbers. (And also assumes that the state space of our universe is uncountably infinite, rather than being countable.) Other work in this vein investigates possible empirically observable effects of “bugs” in the code being run by the simulating computer.
Yet others have considered issues concerning the simulation hypothesis that might charitably be characterized as “philosophical” in nature [32, 21, 26, 37], at least in that they do not involve mathematical analysis of any sort.
Note though that the simulation hypothesis explicitly asks whether one universe (e.g., ours) might be a computer simulation being run in some (perhaps different) universe. Strangely, despite its being formulated this way, there is almost no previous work investigating the simulation hypothesis using the tools of computer science (CS) theory, or related fields like logic theory.
One earlier set of semi-formal results along these lines was presented in [65]. Some of those semi-formal results define “simulation” in terms of the relationship between languages at different levels of the time, polynomial, or exponential hierarchies of CS theory, and some of them define “simulation” in terms of the relationship between languages at different levels the arithmetic or analytic hierarchies of logic, or similar constructions. (See [22, 5, 30] for background on such hierarchies.) Other results in that paper define “simulation” in terms of the relationship between computational machines with different Turing degrees [58, 55]. Yet others consider the application of Gödel’s second incompleteness theorem, defining simulation in terms of languages (in the logic theory sense) with nested sets of axioms.
None of these earlier results in the literature consider what it means for a computational machine to simulate a physical universe, per se. There is no concern for “coupling” the mathematics of CS theory to the laws of physics of our universe. One natural way to address this lacuna is to use a particular formalization of the physical Church-Turing thesis (PCT), and a closely related thesis which I call the reverse PCT (RPCT). That is the approach I adopt in this paper. [^3] The reader unfamiliar with TMs and associated concepts like Universal Turing machines (UTMs), instantaneous descriptions (ID) of the state of a TM and its tapes, prefix-free encodings, etc., should consult Appendix A for appropriate background. Specifically, in this paper I will say that the PCT applies to a particular (portion of a) universe if the dynamics of that (portion of a) universe can be implemented on a Turing machine (TM). The RPCT instead says that any desired TM can be implemented by the dynamics of the universe in question, by appropriate choice of the initial conditions of that universe.
@@ 1.2 General comments
It is important to emphasize that in this paper I do not assume that the PCT applies to our actual physical universe. I do not even restrict attention to those universes that obey the laws of physics of our actual universe, as we currently understand those laws. The focus is instead more general, considering what CS theory has to say about any universe containing a computer that runs a simulation of a universe. In particular, in this paper I am only concerned with establishing the logical possibility of a physical system $V$ that can (contain a computer that can) simulate some universe $V^{\prime}$. I do not investigate the possibility of such a system under the laws of physics as currently understood — never mind the even more narrowly defined question of whether there is such a physical system in our universe, evolving in a manner consistent with the laws of physics as currently understood.
@@
It is also important to emphasize that although I will frequently refer to the computer in a universe as a (U)TM, I do not mean that it is a physical system consisting of a set of infinite tapes with associated heads, etc. Rather I just mean that it has the properties of a (U)TM, i.e., that it is computationally universal. I then choose to discuss this computational system as though it were implemented as a (U)TM. So for example, the universe could be a laptop with a memory that can be extended dynamically an arbitrary finite amount an arbitrary number of times as it runs. (See Section 2.6 below for a detailed example of how a subset of the specific universe occupied by us humans fits into this framework.)
Furthermore, there is no semantics in this paper, only syntax (in the terminology of the foundations of mathematics). There are no structure functions, models, etc. Concretely, this means that there is no distinction between universes that are “real” and those that are “only simulations”. This reflects the viewpoint of the simulation hypothesis itself. Formally, in this paper this lack of explicitly distinguishing real from non-real universes is possible due to my exploiting the physical Church-Turing thesis.
@@ 1.3 Contributions
I begin in Section 2 by presenting the mathematical framework I will use in this paper. I then use that to framework to formally define “simulate” in a way that can apply to any pair of dynamical systems, with no restrictions to the laws of our particular universe. Next, I formally define the PCT in terms of the mathematical framework, as well as the RPCT. These definitions provide the first formalization of the simulation hypothesis, as well as the the first fully generalization formalization of the PCT, applicable to arbitrary universes, not just ours. These definitions also provide the first fully general distinction of the PCT from the RPCT. To help the reader ground the discussion, I also sketch a way to relate the “computers” considered in these definitions to physical subsystems of our actual universe.
@@
In the next section I explicitly prove that if a universe $V^{\prime}$ obeys the PCT, and a universe $V$ obeys the RPCT, then $V$ can simulate $V^{\prime}$. I end that section by describing several arguments based on this result that one might suppose disprove the possibility of self-simulation, i.e., which prove that we could not be simulations in a computer that we ourselves run.
@@
In Section 4 I use Kleene’s second recursion theorem to address these arguments. [^4] This theorem is just called “the recursion theorem” in CS theory; see Appendix B for a summary. I use that theorem to prove that in fact we could be simulations in a computer that we ourselves run. Specifically, I show that if a universe $V$ obeys both the PCT and the RPCT, then it can simulate itself, according to the formal definition of “simulate” provided in Section 2.2. I call this the self-simulation lemma. I then describe several important formal features of the self-simulation lemma, and present an example of how self-simulation might arise with advanced versions of our current laptops. I end this section by describing how self-simulation is a far deeper connection between an entity and itself than arises in all the earlier versions of self-reference considered in the mathematics literature.
@@
In the following section I present several mathematical properties of the number of iterations taken to simulate one’s one dynamics a given time $\Delta t$ into the future. Some of these involve requirements that the time taken to simulate $\Delta t$ into one’s own future does not decrease with $\Delta t$, for any specific pair of values of $\Delta t$. Other properties involve the time-complexity of self-simulation, i.e., how much longer than a time $\Delta t$ it takes to simulate a universe’s evolution up to a time $\Delta t$ in the future.
@@
The next section, Section 6 starts with a a discussion of some of the peculiar philosophical implications of the simulation lemma, and especially of the self-simulation lemma, for notions of identity. In particular, that section contains a discussion of the fact that self-simulation does not just mean that you create some doppelganger of yourself, a clone of yourself, which has autonomy and starts to evolve differently from you once it has been created. Self-simulation does not mean something akin to your stepping into a variant of the Star Trek transporter which creates a copy of you at some other location while the original you still exists. Rather than such cloning of yourself, self-simulation means that you run a program on a computer which implements the exact same dynamics as your entire universe, the universe that contains both you and your computer. So in particular, that universe being simulated in a program running in your computer $N$ contains an instance of you who, in this simulation, is running a program on a computer $N^{\prime}$ that simulates your entire universe, and so in particular simulates an instance of you who, in this simulation-within-a-simulation, is running a program on a computer $N^{\prime\prime}$ that simulates your entire universe, and so in particular … Crucially, under the PCT, all those instances of you are you; it is meaningless to ask which of those instances “is the real you”, with the others being “just a copy”. This section ends with a discussion of the philosophical quirks that would arise if the program being used to (self-)simulate universe is encrypted, so that only the being with a special decryption key can understand the result of that computation.
@@
The simulation and self-simulation lemmas allow us to define the “simulation graph”. This is the directed graph where each node is a universe containing a computer, and there is an edge from one node to another if the (universe identified with the) first node can simulate the (universe identified with the) second node. In Section 7 I present a preliminary investigation of this graph.

### v3

# What computer science has to say about the simulation hypothesis

@@ Abstract
The simulation hypothesis has recently excited renewed interest in the physics and philosophy communities. However, the hypothesis specifically concerns computers that simulate physical universes. So to formally investigate the hypothesis, we need to understand it in terms of computer science (CS) theory. In addition we need a formal way to couple CS theory with physics. Here I couple those fields by using the physical Church-Turing thesis. This allow me to exploit Kleene’s second recursion, to prove that not only is it possible for us to be a simulation being run on a computer, but that we might be in a simulation being run a computer by us. In such a “self-simulation”, there would be two identical instances of us, both equally “real”. I then use Rice’s theorem to derive impossibility results concerning simulation and self-simulation; derive implications for (self-)simulation if we are being simulated in a program using fully homomorphic encryption; and briefly investigate the graphical structure of universes simulating other universes which contain computers running their own simulations. I end by describing some of the possible avenues for future research. While motivated in terms of the simulation hypothesis, the results in this paper are direct consequences of the Church-Turing thesis. So they apply far more broadly than the simulation hypothesis.
@@ 1.1 Background
It is broadly supposed in the physics community that the future values of any properties that we can experimentally measure concerning a system in our universe can be computed ahead of time on a Turing machine [56, 54]. Loosely speaking the supposition is that your laptop, sufficiently “souped up”, could be used to simulate the future value of any physical variable you could experimentally measure, if you just fed it sufficient information about the current state of the universe. Concretely, there has never been an experiment proposed where we would fully specify the initial condition of a closed system, would then measure the future value of a property of that system, and yet for some reason or other it was theoretically impossible for us to compute the future value of that property ahead of time on a sufficiently souped-up laptop.  [^1] Throughout this paper, I will often refer to a “Turing machine” as this kind of device, defined in terms classical physics, even though the physical universe is quantum mechanical. There is a large literature on the “quantum Church Turing thesis”, applicable to quantum computers. The distinction between these two type of TM and associated form of the PCT are tangential to the focus of this paper, and therefore I will not engage with those distinctions. See Appendix A
Following much of the literature, I will refer to this supposition as the “physical Church-Turing thesis’ (PCT)’ ([56, 54] and references therein). [^2] The precise formulation of the PCT is the subject of much debate. Here, for reasons of expediency, I simply define it to mean the supposition described just above. The PCT is particularly compelling if one adopts the “multiverse” perspective common in physics, which says that our universe is a mathematical structure, given by computing the logical consequences of a set of fundamental mathematical laws [63, 64, 34, 44, 41].
Under the PCT, it is impossible for us to experimentally rule out the possibility that the evolution of all properties of our physical universe that we can measure are the outputs of a program being run on some Turing machine (or even more powerful computational machine). [^3] See Appendix A for formal definitions of Turing machines and their properties. More loosely, the PCT is consistent with the hypothesis that our universe is a simulation being run on some Turing machine [21]. This second possibility is known in the literature as the “simulation hypothesis”. In this paper, to avoid ontological issues, I will use the term “simulation hypothesis” more broadly, to refer to either of these two possibilities, without distinguishing them.
In addition to the PCT, another common supposition in the physics community is that our physical universe can contain a subsystem whose dynamics implements an arbitrary Turing machine (TM) [40]. In other words, the default stance in much of the physics community is not to rule out the possibility that a given hypothesized dynamical system can exist in our universe, simply because that system has the computational power of a TM.  [^4] Indeed, there has been some very important work that considers whether determining certain properties of physical systems might require computational power that is equal to that of a (universal) TM [26, 59]. Crucially, these properties are not experimentally measurable though, which is why they do not contradict the PCT.
In contrast to the PCT, I will refer to this second supposition as the “reverse physical Church-Turing thesis (RPCT)”. I will also sometimes use the shorthand “CT thesis” to refer to the two suppositions at once. [^5] The reader unfamiliar with TMs and associated concepts like Universal Turing machines (UTMs), instantaneous descriptions (ID) of the state of a TM and its tapes, prefix-free encodings, etc., should consult Appendix A for appropriate background.
Note that the RPCT raises the possibility that whether or not our universe is a simulation, our universe could contain within it a (physical subsystem implementing a) computer which is running a simulation of some other universe. Intriguingly, it would seem that combining this implication of the RPCT with the PCT would establish that our universe can contain a computer that is implementing a simulation of our universe itself.
Whether or not our universe does actually obey the PCT and / or the RPCT, there are many questions that they raise;
  1. 1.
    What subsystems in our physical universe validate the RPCT (if any)?
  2. 2.
    As a practical matter, assuming there are subsystems of our universe that validate the RPCT, how much of their computational power is accessible to us humans?
  3. 3.
    What are the implications of computer science (CS) theory for the limitations of any physical universe obeying the full CT thesis?
  4. 4.
    In particular, what are the implications of CS theory for the simulation hypothesis?
The first two questions in this list have previously been considered in the physics literature, relatively extensively [47, 46, 15, 26, 59]. There has also been work in the physics community that addresses the third question in the list [63, 64, 10, aaro05, 34]. In addition there has been some work in the physics community that focuses on the fourth question, investigating how we might experimentally determine whether the simulation hypothesis holds [14, 11, 9]. [^6] In addition to this work in the physics community, there has been a lot of work in the philosophy of science literature on the simulation hypothesis. See Appendix E, for a critical summary of this work, in particular discussing the problematic aspects of the attempts in that literature to assign a probability to the simulation hypothesis.
However, with one partial exception discussed in Appendix F, no work has even been done on using CS theory to investigate the simulation hypothesis. This is despite the fact that, after all, the simulation hypothesis concerns computer simulation. So one would expect that the very powerful results of CS theory are extremely important to understanding the simulation hypothesis. Phrased differently, if we don’t investigate what CS theory has to say about fourth question, it seems likely that deeply important aspects of the simulation hypothesis will be missed.
One natural way to address this hole in the literature starts by formalizing both the PCT and the RPCT in a very general (and fully rigorous) way. This would allow us both to formalize what it would mean for our physical universe to be simulated, and what it would mean for some subsystem in our universe to simulate a universe. Armed with this machinery, we could then apply the powerful CS-theoretic results concerning TMs, to investigate question four. I adopt this approach in this paper.
As an example of the benefits of this approach, I show below how Kleene’s second recursion theorem from CS theory can be used to formally investigate this possibility, that our physical universe could in theory simulate itself. On the other hand though, below I will also use a different theorem from CS theory, Rice’s theorem, to derive impossibility results concerning simulation of one universe by another. (The reader unfamiliar with one or the other of these two theorems should read Appendix B.)
It is important to emphasize that the “simulation hypothesis” as popularly conceived is that there is some being running a computer program in some physical universe, and all observables in our universe are just a set variables in that program. What the program does is evolve those variables, from one time to another. In particular, in this popular conception of the simulation hypothesis, we are just such variables in a computer program, as are all our experimental apparatuses and observations.
While motivated in terms of that popular version of the simulation hypothesis, the actual results in this paper are far more general. In fact, they don’t even require that there be such a being who is running a program on a computer that may be simulating some universe..
Rather, the analysis in this paper starts from the fact that all physical experiments are interactions between an experimenter and physical reality. If we assume the Church-Turing thesis, the responses of physical reality during such an interaction can be modeled as the results of a computation on a TM. The Church-Turing thesis also means that the behavior of we the experimenters can be modeled that way. One might also assume that we ourselves can run a Turing complete computer. This paper is simply an investigation of the formal consequences if adopt one and / or the other of these two assumptions. Those formal consequences would apply to our actual universe if those two assumptions hold, i.e., they are a new set of (im)possiblitiy results concerning the laws of physics.
@@ 1.2 General comments
It is important to emphasize that in this paper I do not assume that the PCT applies to our actual physical universe, and / or that the RPCT does. Nor do I make any assumptions about whether the simulation hypothesis does or does not hold for our universe. These issues are ultimately decided by the Hamiltonian of our actual universe. However, I do not even restrict attention to those universes that obey the laws of physics of our universe, as we currently understand those laws. The focus is instead more general, considering what CS theory has to say about any universe whose dynamics could (in theory at least) be simulated on a (Turing complete) computer, and / or a universe that contains a computer that could simulate the dynamics of a universe. (See [14, 11, 9] for some work in the physics community that touches on the issue of how we might experimentally determine whether the simulation hypothesis holds.)
@@
In particular, in this paper I am only concerned with establishing the mathematically necessary properties of a universe $V$ that can (contain a computer that can) simulate some universe $V^{\prime}$. I do not investigate the physical possibility of such a computer under the laws of physics in our universe, as we currently understand those laws. Nor do I consider the even more narrow question of whether there is such a computer in our universe.
@@
It is also important to emphasize that I will frequently refer to the computer in a universe as a universal Turing machine (UTM). By this I do not mean that it is a physical system consisting of a set of infinite tapes with associated heads, etc. Rather I just mean that it has the properties of a (U)TM, i.e., that it is computationally universal. I then choose to discuss any such computational system as though it were implemented as a (U)TM. So for example, the universe could be a running laptop with a memory that can be extended dynamically by an arbitrary finite amount, for an arbitrary number of times. This is the standard approach in the literature for justifying the application of Turing machine theory to our physical universe. (See Section 2.6 below for a detailed example of how a subset of the specific universe occupied by us humans fits into this framework.)
Furthermore, to adopt the terminology of logic and model theory, there is no semantics considered in this paper, in the sense of T-schema, etc. So there are no structures, definable relations, etc.. This reflects the implicit viewpoint of the simulation hypothesis itself. Formally, just like the simulation hypothesis itself, my use of the the physical Church-Turing thesis can be viewed as coupling syntax with semantics, but I have no more explicit consideration of semantics. (See [63, 64, 10, aaro05] for work in the physics literature that couples syntax and semantics in a similar way, and the discussion of ontic structural realism in Appendix E for related, informal discussion.) Just as in the simulation hypothesis, I make no distinction between universes that are “real” and those that are “only simulations”.
As a final comment, the analysis in this paper only considers deterministic TMs. The simulator and simulatee are formalized as such TMs. However, many scenarios in CS theory involve probabilistic TMs, e.g., the complexity class BPP, interactive proof systems, cryptography, etc. In addition, one might argue that to extend the analysis in this paper to a quantum mechanical universe, we would need to allow for probabilistic dynamics in the simulated universe, e.g., if one adopts a Copenhagen interpretation of quantum mechanics.
One can easily extend the formalism in this paper to handle such stochasticity in the simulated TM and / or the simulator’s TM. To do so we just need to suppose that the deterministic TMs have an auxiliary, read-only input tape they can use, which is filled with an infinite string of random bits in addition to their other inputs. (This is the standard trick in CS theory for defining TMs that behave IID randomly even though they are purely deterministic TMs. See [5].)
@@ 1.3 Contributions
The focus of this paper is on the CS theoretic aspects of the simulation hypothesis.
@@
I begin Section 2 by presenting the mathematical framework I will use in this paper. The simulation hypothesis in particular, and both the PCT and RPCT more generally, all involve a pair of dynamical systems, e.g., the simulating universe and the simulated universe. Accordingly, I next formally define “simulate” in a way that can apply to any pair of dynamical systems, with no restrictions to the laws of our particular universe. After that I formally define the PCT in terms of the mathematical framework I will use in this paper, as well as the RPCT. These definitions allow me to provide the first formalization of the simulation hypothesis, as well as the first fully general formalization of the PCT, applicable to arbitrary universes, not just ours. These definitions also provide the first fully general distinction of the PCT from the RPCT. For those readers who have not previously considering the PCT and / or RPCT in detail, I end this section by giving some pointers to the literature on physical systems that are computationally universal. I also introduce a new example of such a scenario, grounded in purely classical physics, and without requiring ability manipulate and observe real numbers ot infinite precision.
@@
In the next section I explicitly prove that if a universe $V^{\prime}$ obeys the PCT, and a universe $V$ obeys the RPCT, then $V$ can simulate $V^{\prime}$. I end that section by describing several semi-formal arguments based on this result that one might suppose disprove the possibility of self-simulation, i.e., which prove that we could not be simulations in a computer that we ourselves run.
@@
In Section 4 I use Kleene’s second recursion theorem to address these semi-formal arguments against self-simulation. [^7] This theorem is just called “the recursion theorem” in CS theory; see Appendix B for a summary. I do this by using that theorem to prove that in fact we could be simulations in a computer that we ourselves run. Specifically, I show that if a universe $V$ obeys both the PCT and the RPCT, then it can simulate itself, according to the formal definition of “simulate” provided in Section 2.2. I call this the self-simulation lemma. I then describe several important formal features of the self-simulation lemma, and present an example of how self-simulation might arise with advanced versions of our current laptops. I end this section by describing how self-simulation is a far deeper connection between an entity and itself than arises in all the earlier versions of self-reference considered in the mathematics literature.
@@
In the following section, Section 5, I present several mathematical properties of the number of iterations taken to simulate one’s dynamics a given time $\Delta t$ into the future. Some of these properties involve the implications if we require that the time taken to simulate $\Delta t$ into one’s own future does not decrease with $\Delta t$, for any specific pair of values of $\Delta t$. Other properties involve the time-complexity of self-simulation, i.e., how much longer than a time $\Delta t$ it takes to simulate a universe’s evolution up to a time $\Delta t$ in the future.
@@
The next section, Section 6 starts with a a discussion of some of the peculiar philosophical implications of the simulation lemma, and especially of the self-simulation lemma, for notions of identity. In particular, that section contains a discussion of the fact that you self-simulating does not just mean that you create some doppelganger of yourself, a clone of yourself, which has autonomy and starts to evolve differently from you once it has been created. Self-simulation does not mean something akin to your stepping into a variant of the Star Trek transporter which creates a copy of you at some other location while the original you still exists.
Rather than such cloning of yourself, self-simulation means that you run a program on a computer which implements the exact same dynamics as your entire universe, the universe that contains both you and your computer. So in particular, that universe being simulated in a program running in your computer $N$ contains an instance of you who, in this simulation, is running a program on a computer $N^{\prime}$ that simulates your entire universe, and so in particular simulates an instance of you who, in this simulation-within-a-simulation, is running a program on a computer $N^{\prime\prime}$ that simulates your entire universe, and so in particular … Crucially, under the PCT, all those instances of you are you; it is meaningless to ask which of those instances “is the real you”, with the others being “just a copy”. [^8] As discussed in Section 5, an instance of your universe that you are simulating would be “time-delayed”, in that as you observed that simulation unfold, after it has run for a certain amount, you would notice that it is running more slowly than your universe. In particular, the simulation unfolding would be the same as at delayed version of you. Of course, there is no experimentally testable way that the simulated version of you could notice this.
This section ends with a discussion of the philosophical quirks that would arise if the program being used to (self-)simulate a universe is encrypted as it runs, as in homomorphic computing [3]. In such situation only the being with a special decryption key can understand the result of that computation. In particular, I raise the issue of what would happen in this situation if that being running the computation – running our universe – were to lose the decryption key, and so no longer could understand or observe any aspect of our universe.
@@
Returning to mathematics, the simulation and self-simulation lemmas allow us to define the “simulation graph”. This is the directed graph where each node is a universe containing a computer, and there is an edge from one node to another if the (universe identified with the) first node can simulate the (universe identified with the) second node. In Section 7 I present a preliminary investigation of this graph.